[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Aug 09, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Aug 10 03:05:01 UTC 2017


HSU Shift Log for Wed Aug 09 2017 23:04:55 GMT-0400 (EDT)

Shift: R

Forecaster: L Avila

Technical Issues Summary:

ATCF issues: ATCF server lost mounted directory access due to  
redhat "rbind" bug. Was unable to transmit 21Z package until Doug Gaer  
manually restarted the service as root. (DAZ)

N-AWIPS issues: 4 N-AWIPS machines lost mounted directory access. Rebooted  
Sawfish which fixed that one. Doug Gaer addressed others. (DAZ)

AWIPS-2/NCP issues: None

Other Techincal issues: Comp server lost mounted directory access for same  
reasons listed earlier. Graphics failed, had to be re-run. GTWO/frontpage  
graphic failed, had to rerun manually from comp server by changing  
timestamp of xml advisory file after Doug Gaer manually fixed the redhat  
issue. Refer to email from Pablo for more details (DAZ) Rat traps placed in  
Room 113 and the FEMA room due to evidence that Rats were eating mangoes  
left in those rooms.

Miscellaneous issues: Chris Mello assisted with 8pm outlooks as an HSM.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170810/7ca47f9e/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list